Grade 11Discuss with Askiitians Tutors

the limiting points of coaxal system of circles defined by x^2+y^2-6x-6y+4=0 & x^2+y^2-2x-4y+3=0 are

reshma lr
12 Years agoGrade 11

1 Answer

Mukesh Sharma
12 Years ago
first let s=x^2+y^2-6x-6y+4=0 ans s1=x^2+y^2-2x-4y+3=0
then find L=sminus s1(s-s1) then find s+ ?L
so we will get equation and separate x,y and c component then we find radius of circle as
R= sqare root xcomponent+y component +c =0 thereafter we get an quadratic eqation we will solve that and hence find values those will be limiting points